solve each system of equations by substitution method
3x+6y=9
4x+8y=16

Answer by ikleyn(52794)

3x + 6y = 9    (1)
4x + 8y = 16   (2)

Let us divide equation (1) by 2 (both sides) and equation (2) by 4 (both sides). You will get

x + y = 3      (1')
x + y = 4      (2')

Note that the left sides of these equations (1') and (2') are identical, while left sides are different.
It means that the system (1')-(2') is a) dependent and b) inconsistent. It has no solution. 
It implies that the original system (1)-(2) is a) dependent and b) inconsistent, too. It also has no solution. 

One can see it from the very beginning that the system (1)-(2) is dependent and inconsistent.
Indeed, the equations (1) and (2) have proportional left side coefficients, while the right sides are proportional with the different coefficient of proportionality.